OpenStudy (pphalke):

soybean meal is 14% protein; commeal is 7% protein. How many pounds of each should be mixed together in order to get 280 LB that is 10% protein?

OpenStudy (shamim):

LB means?

OpenStudy (scooby6363):

@shamim LB = Pounds

OpenStudy (shamim):

Ok

OpenStudy (shamim):

Let soybean=x

OpenStudy (shamim):

Commeal=280-x

OpenStudy (shamim):

U know 14%=0.14

OpenStudy (shamim):

7%=0.07

OpenStudy (pphalke):

x+y=280 0.14x+0.01y=(280)(0.07) so far good?

OpenStudy (shamim):

Total amount of protein in soybean=0.14*x

OpenStudy (pphalke):

check my above which says 0.14x+0.1y=(280)(0.07)

OpenStudy (shamim):

0.14*x+(280-x)*0.07=280*0.1

OpenStudy (shamim):

x=?

OpenStudy (pphalke):

no idea

OpenStudy (shamim):

Wt do u think abt my equation

OpenStudy (pphalke):

@Scooby6363

OpenStudy (shamim):

Can i explain my equation?

OpenStudy (pphalke):

yes please explain @jim_thompson5910

OpenStudy (shamim):

Totsl amount of protein in soybean=0.14*x Right?

OpenStudy (pphalke):

yes

OpenStudy (shamim):

Total amount of protein in commeal=0.07*(280-x)

OpenStudy (pphalke):

yes

OpenStudy (shamim):

So sum of protein in soybean + commeal=0.14x+(280-x)*0.07

OpenStudy (shamim):

Right?

OpenStudy (pphalke):

yes

OpenStudy (shamim):

If protein is 10%

OpenStudy (shamim):

280 pound had total protein=280*0.1=?

OpenStudy (pphalke):

its 28

OpenStudy (shamim):

Ok

OpenStudy (shamim):

So we can write 0.14*x+(280-x)*0.07=28

OpenStudy (shamim):

x=?

OpenStudy (pphalke):

no idea

OpenStudy (shamim):

0.14x+280*0.07-0.07*x=28

OpenStudy (shamim):

Right?

OpenStudy (pphalke):

yes

OpenStudy (shamim):

After multiplicatikn we get my lat equation

OpenStudy (shamim):

280*0.07=?

OpenStudy (pphalke):

19.6?

OpenStudy (shamim):

Ok

OpenStudy (pphalke):

did u read the question?

OpenStudy (shamim):

0.14x-0.07x=?

OpenStudy (pphalke):

0.07

OpenStudy (shamim):

Now after very few calculations tell me x=?

OpenStudy (pphalke):

how would i tell you that do i do something to 1.96 and 0.07 and di u read the question???

OpenStudy (shamim):

Can we write 0.07x=28-19.6 Right?

OpenStudy (pphalke):

x=372

OpenStudy (shamim):

Can we write 0.07x=28-19.6 Right?

OpenStudy (pphalke):

yes 0.07x=26.04 then x=372 how does that get me to the answer

OpenStudy (pphalke):

@shamim

OpenStudy (shamim):

28-19.6=8.4 Right?

OpenStudy (pphalke):

no its 26.04

OpenStudy (shamim):

U r wrong!!

OpenStudy (shamim):

Tell me 28-19=?

OpenStudy (pphalke):

thats 9

OpenStudy (shamim):

Now tell me 28-20=

OpenStudy (shamim):

Now tell me 28-19.6=?

OpenStudy (pphalke):

no no @shamim i used my calculator its 28-1.96 and that came out 26.04 check ur calculations

OpenStudy (pphalke):

oh ok yes i see my error

OpenStudy (pphalke):

then x=120

OpenStudy (shamim):

It will b 28-19.6=?

OpenStudy (shamim):

So amount of soybean=120 pound

OpenStudy (pphalke):

what about amount of cornmeal?????

OpenStudy (shamim):

Amount of commeal=?

OpenStudy (pphalke):

i dont know

OpenStudy (shamim):

Amount of commeal=280-x=280-120=?

OpenStudy (pphalke):

160

OpenStudy (shamim):

Correct!!

OpenStudy (shamim):

Did u get it all ?